OpenStudy (anonymous):

-4y + 17 < or equal to 29 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (anonymous):

-4y + 17 <= 29 -4y<=12 -17 y>=-3 divide by-4 so the < sign flips because of division by a negative

OpenStudy (anonymous):

I hink you mus have forgotten an x somewhere @analiza but if you didn't then y>=-3 is correct and looks like